OpenStudy (anonymous):

Solve for the unknown side in each right triangle to the nearest tenth. C=8 A=6 B=? What is the formula to find out what B is?

OpenStudy (anonymous):

a^2 + b^2 = c^2

OpenStudy (anonymous):

in right triangles you can use pythagorean theorem a^2+b^2 = c^2 so 6^2-(8^2) = b^2

OpenStudy (anonymous):

Thaaaaanks!

OpenStudy (anonymous):

you are welcome!

OpenStudy (anonymous):

no problem!